OpenStudy (anonymous):

find this in simplest form > On a hike, Peter and Jamie walked 31/8 hours at 21/2 miles per hour. How far did they walk? for exsample 7/5 aka 7 over 5 over 11 >

OpenStudy (texaschic101):

Distance = speed x time D = 31/8 x 21/2 D = 651/16 miles ??? that just looks weird....you might want to get a second opinion

OpenStudy (texaschic101):

651/16 = 40 11/16 miles